5. At a conference in 2006, an executive from a brokerage firm in the money market, told a group of analysts that 70% of investors were confident in achieving their investment objectives. USB investor conducted a study with information from 300 investors and found that 67% were confident in achieving their investment objectives. Should the executive's claim be rejected, at 95% confidence? Interpret the result
Answer #1

Solution:

p0 = 70% = 0.70

Claim : p = 0.70 or not

The null and alternative hypothesis can be written as

H0 : p = 0.70 vs H1 : p 0.70

sign in H1 indicates the TWO tailed test.

n = 300

Let denotes the sample proportion

= 67% = 0.67

We test the claim using 95% confidence interval.

c = 0.95

= 1- c = 1- 0.95 = 0.05

  /2 = 0.05 2 = 0.025 and 1- /2 = 0.975

Search the probability 0.975 in the Z table and see corresponding z value

  = 1.96

Now , the margin of error is given by

E = /2 *  

= 1.96 * [(0.67* (1 - 0.67)/300]

=  0.0532

Now the confidence interval is given by

( - E)   ( + E)

(0.67 - 0.0532)   (0.67 - 0.0532)

0.6168 0.7232

Interval is (0.6168 , 0.7232 )

Observe that , the hypothetical value 0.70 is in the interval (0.6168 , 0.7232)

So, we fail to reject the null hypothesis H0

We can there is no evidence to reject the claim that the proportion of investors that were confident in achieving their investment objectives is 70%.

i.e. The proportion of investors that were confident in achieving their investment objectives is not different from 70%.